40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12616     Entry time: Tue Nov 15 19:22:17 2016
Author: gautam 
Type: Update 
Category: General 
Subject: housekeeping 

PRM and SRM sat. boxes have been switched for some time now - but the PRM sat. box has one channel with a different transimpedance gain, and the damping loops for the PRM and SRM were not systematically adjusted to take this into account (I just tweaked the gain for the PRM and SRM side damping loops till the optic damped). Since both sat. boxes are nominally functioning now, I saw no reason to maintain this switched configuration so I swapped the boxes back, and restored the damping settings to their values from March 29 2016, well before either of this summer's vents. In addition, I want to collect some data to analyze the sat. box noise performance so I am leaving the SRM sat. box connected to the DAQ, but with the tester box connected to where the vacuum feedthroughs would normally go (so SRM has no actuation right now). I will collect a few hours of data and revert later tonight for locking activities....

ELOG V3.1.3-